OpenStudy (anonymous):

Find the number which when added to the terms of the ratio 19:17, makes it equal to the ratio 11:10 shuld be added to ratio ans-- a-1, b-2, c-3, d-4

OpenStudy (anonymous):

C

OpenStudy (anonymous):

how ?

OpenStudy (anonymous):

19+3=22 17+3=20 22:20=11:10

OpenStudy (anonymous):

kkk !!

OpenStudy (anonymous):

(19+x)/(17+x)=11/10 190+10x=11x+187 x=3